OpenStudy (anonymous):

Integrate (e^2x)/(1 + e^x) @Calculus1

OpenStudy (anonymous):

u sub it via \[u=e^x, du = e^xdx\] to get \[\int \frac{u}{u+1}du\]

OpenStudy (turingtest):

what happened to the two sat?

OpenStudy (anonymous):

oh \[e^x\times e^xdx=e^{2x}\] in the numerator

OpenStudy (anonymous):

sorry i meant \[e^{2x}dx\]

OpenStudy (anonymous):

that is why it is just \[\int\frac{udu}{u+1}\]

OpenStudy (anonymous):

still have to integrate though. since power is the same top and bottom divide to get \[\frac{u}{u+1}=1-\frac{1}{u+1}\] then integrate to get \[u-\ln(u+1)\] and finally \[e^x-\ln(e^x+1)\]

OpenStudy (turingtest):

+C !!!!!!!!

OpenStudy (anonymous):

I'm still not understanding why the 2 is irrelevant...

OpenStudy (turingtest):

He integrated the form above, so it is relevant. If it was irrelevant you would integrate without a u in the denominator, just a du. because he wrote \[u=e^x \to du=e^xdx\]the numerator is now \[udu=e^{2x}dx\]Then he did his fraction trick to the expression and was able to integrate it.

OpenStudy (anonymous):

Can you show a bit more of how the fraction was split?

OpenStudy (turingtest):

I was just wondering and figure it out... clever\[{u \over u+1}={u+1-1\over u+1}={u+1\over u+1}-{1\over u+1}=1-{1\over u+1}\]That's why satellite gets props from me...

OpenStudy (anonymous):

oh thank you for the compliment. but in fact i did not do that trick. i just divided

OpenStudy (turingtest):

how did you "just divide" without a 1 in the numerator? I've never done that.

OpenStudy (anonymous):

i will see if i can write it it is one step

OpenStudy (anonymous):

ok i cant write division here but it does like this basically you divide u by u + 1 u goes in to u one time 1 times u + 1 is u + 1 subtract and get -1 that is the remainder. so answer is \[1-\frac{1}{u+1}\] i can never figure out how to write a long division here

OpenStudy (anonymous):

Thanks for the lessons sat and turing!

OpenStudy (turingtest):

Oh I see, that makes sense, I'm surprised I've never seen it before. I guess you could draw itThanks!!!!!!! :D